# SIGGesture: Generalized Co-Speech Gesture Synthesis via Semantic Injection with Large-Scale Pre-Training Diffusion Models

### v1

@@ 3.2 Semantic Gesture Collection
Figure 2: The statistical data of semantic gesture dataset (upper left part), some examples of semantic label (lower left part), and some semantic gesture examples (right part).
@@
All of semantic gesture clips are encoded into latent embedding $V=\{v_{0},v_{1},...,v_{N}\}$ by a pre-trained VQVAE model. To improve the diversity of semantic gesture, we add a random noise to the embeddings during semantic injection.
3.3 Diffusion models with semantic injection
@@
SIGGesture employs a diffusion model to synthesize the latent representation of gesture, which consists of a diffusion process and a denoising process, as shown in upper part of Figure 1. Given a specific real motion data distribution $p(x_{0})$, our goal is learn a distribution $q_{\theta}(x_{0})$ that is parameterized by ${\theta}$.
@@
Diffusion process. We follow the previous DDPM [23] definition of diffusion as a Markov noising process with latents $x_{0:T}$ that follow a forward noising process $p(x_{0:T}|x_{0})$. Here, $x_{0}$ is the latent motion data and same as previous notation $\hat{e}$. The forward diffusion process is defined as a Markov chain that gradually adds Gaussian noise to the latent representation of motion, as follows
$\begin{split}q(x_{0:T}|x_{0})=\prod_{t=1}^{T}q(x_{t}|x_{t-1}),\end{split}$ (1)
@@
$\begin{split}q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}{I}).\end{split}$ (2)
$\beta_{t}$ are variance schedule hyperparameters that controls the distribution of Gaussian noise. In the reverse denoising process, $\beta_{t}$ are pre-defined constant parameters. By this diffusion process, the original latent representation of motion $x_{0}$ progressively substituted by random noises. When $T$ approaches infinity, the distribution of $x_{T}$ follows a pure white noise distribution.
Denoising process. Through diffusion process, the real data are transferred into pure white noise by adding random noise. On the contrary, the denoising process recovers the real data from pure white noise by removing the random noise. The unconditional reverse process of recover a gesture motion can be represented by the following formula
$\begin{split}p_{\theta}(x_{0:T})=p_{\theta}(x_{0})\prod_{t=1}^{T}p_{\theta}(x_{t-1}|x_{t}),\end{split}$ (3)
@@
$\begin{split}p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t};u_{\theta}(x_{t},t),\beta_{t}I).\end{split}$ (4)
The noise $x_{t}$ is sampled from the following defined distribution
$\begin{split}p(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha_{t}}},(1-\bar{\alpha_{t}}){I}).\end{split}$ (5)
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha_{t}}=\prod_{k=0}^{t}\alpha_{k}$. The above explanation shows the process of unconditioned diffusion generation. For conditional content synthesis, the conditions should be injected into the diffusion generation procedure. The details of the denoising network are shown in the upper right part of Figure 1. In our method, the conditions are audio $c_{audio}$ and speaker identity $c_{id}$ respectively. For simple notation, we mark the two conditions as $c=[c_{audio},c_{id}]$. It should be noted that the audio feature is extracted by Jukebox [13] instead of WavLM [12]. Jukebox [13], trained on large-scale music datasets, can focus on the rhythmic features rather than the speech semantic content. By injecting the condition $c$ into the generation process, the reverse process of each time step can be updated by the following formula
$\begin{split}p_{\theta}(x_{t-1}|x_{t},c)=\mathcal{N}(x_{t};u_{\theta}(x_{t},t,c),\beta_{t}I).\end{split}$ (6)
@@
Training objective. For a speech condition $c$, we reverse the forward diffusion process by learning to estimate $p(x_{t},t,c)$ approximate $x_{0}$ with a parameterized model for all times step $t$. We use a loss function proposed in [23] to optimize the whole model, as following,
$\begin{split}\mathcal{L}={E_{x,t}[\parallel x-p(x_{t},t,c)\parallel^{2}_{2}]}.\end{split}$ (7)
@@
During training, the model is trained under conditional and unconditional manner with a specific probability. Once the training converges, the the diffusion model can predict the noise parameters by considering the unconditional setting and condition $c$.
@@
Inference with semantic injection. The inference process involves solely the denoising process of diffusion models. The pipeline of inference process is shown in the lower part of Figure 1. To be specific, speech with its corresponding text will be feed into the generation pipeline. It should be noted that the words in the text is aligned with speech in timeline. Each words have a duration time. Then, the text are fed into the LLMs to produce candidate semantic gestures $\mathcal{G}=(G_{0},...,G_{t})$ for the speech content. The interaction with LLMs process is shown in the left-top corner of inference pipeline in Figure 1. With the ability of LLM, the proposed method can process various language, such as Chinese and Japanese.
@@
The candidate semantic gesture is encoded into latent embedding $\mathcal{P}=(p_{0},...,p_{n})$, where $n$ is number of candidate semantic gestures. We set a control $K$ to determine the degree of semantic injection. From time step $T$ to $K$, the denoising latent variable is marked as $\hat{x_{t}}$ that contains the semantic information.
$\begin{split}p_{\theta}(x_{K:T})=p_{\theta}(x_{T})\prod_{t=K}^{T}p_{\theta}(x_{t-1}|\hat{x}_{t}).\end{split}$ (8)
@@
The semantic gesture $P$ can be injected into the latent variable $x_{t}$ by the follow formula
$\begin{split}\hat{x}_{t}=x_{t}*m+P*(1-m),\end{split}$ (9)
@@
where $m$ denotes the timeline mask of semantic words. From time step $K$ to 1, the denoising latent variable $x_{t}$ is not with semantic injection, which can preserve the structure and smoothness of the generated gesture.
@@
Generating long sequence. For synthesizing long motion sequence, the common practice is autoregressive method, which generates content sequentially. However, generative models trained on small pieces may cause unnaturalness and defectiveness. Instead of auto-regressive manner, we adopt DiffCollage [51] to generate long motion sequence.
3.4 Large-scale motion corpus collection
@@
In 3D animation, high-quality data collection is remarkably expensive, professional, and time-consuming. High-quality motion capture requires specialized cameras, suits, and markers. Besides, it often requires experienced technicians, actors, and animators to set up, perform, and process the captured data. The process of capturing, cleaning, and integrating motion data is labor-intensive and time-consuming, contributing to the overall expense. In recent years, several high-quality datasets are released to the research community, as shown in Table 1. BEAT [32] is the largest dataset for co-speech gesture synthesis, which contains 76 hours data. However, data-driven approaches with larger models are also short for data. In the inspiration of large-language training paradigm, noise data for pre-training and high-quality data for fine-tuning, we aim at collecting a large-scale gesture corpus for pre-training by video motion capture technique such as HybridIK [29], SMPLify [5]. Before running the video motion capture approach, we collect a large mount of videos about speech or talk from the Internet, such as TED. Then, we conduct the data cleaning pipeline to obtain the final shots of interests. Table 2 shows the statistical data of the motion corpus collection. Finally, we obtain about 400 hours 3D gesture assets of speech. As shown in Table 1, the proposed dataset is significantly larger than previous datasets by at least 6 times.
Table 2: Statistical data of the motion corpus collection.
@@
4 Experiments
4.1 Datasets
In this paper, we use the largest high-quality speech-gesture dataset BEAT to evaluate the proposed method. BEAT is constructed by using commercial motion capture system, including face emotion and body motion. It contains about 76 hours motion-audio paired data of 30 speakers talking about different topics. We deeply clean the dataset by removing some defective motion sequences such as speaker drinking, wandering, and incorrect motion. After cleaning the raw data, the BEAT dataset contains 46 hours motion data and its corresponding text and audio. During training the VQVAE model, the original Euler angles are converted to rotation matrix for better convergence. During training the diffusion-based generation model, all motions are represented by latent embeddings in VQVAE latent space. The datasets collected from Internet by video motion capture are only used for pre-training because of its relatively poor quality.
4.2 Evaluation Metrics
For content generation tasks, human objective evaluation is the most important evaluation method because of lacking clear ground and explicit criteria.Therefore, the human objective evaluation is the main evaluation method in our method. Each motion slice is re-targeted to Mixiamo model and rendered as a motion video sequence. They are asked to rate the slices from the following four aspects respectively: (1) Naturalness, (2) Rhythm, (3) Diversity, (4) Semantic. The scores assigned to each rating are in the range of 1-10, corresponding from worst to best. All scores are normalized to 0-1. Besides, we also calculate some quantitative metrics adopted by previous works. FGD [49] measures the distribution difference between generated data and ground truth. Beat Consistency (BC) [30] calculates the average distance between every audio beat and its nearest motion beat. Diversity is a metric of measuring the variations of the synthesized gesture. Semantic-Relevant Gesture Recall (SRGR) [32] measures the semantic correct key point of synthesized data by comparing it with the ground truth data.
4.3 Implementation Details
The whole training process consists of two steps. In the first step, we train the VQVAE model using the Adam optimizer for 400 epochs. In the second step, we train the diffusion-based gesture generation model using the Adam optimizer for 3000 epochs with batchsize 256. The diffusion steps for training and inference are 1000. During the training of diffusion model, we firstly use the collected data for pre-training (2000 epochs with batchsize 256), and then use the high-quality dataset for fine-tuning (1000 epochs). The length of the input audio feature is 12 seconds. We train the diffusion model on 32 V100 GPUs.
4.4 Comparing Baselines
We compare our method on BEAT dataset with the state-of-the-art methods in the last year, including DiffuseStyleGesture[44], TalkSHOW[48], QPGesture [45], EMAGE [31]. DiffuseStyleGesture[44] is a diffusion model-based co-speech gesture synthesis approach, which takes the motion style into generation process. TalkSHOW[48] uses a simple full-body speech-to-motion generation framework with auto-regressive manner, in which the face, body, and hands are modeled separately. QPGesture [45] is a quantization-based and phase-guided motion-matching approach for co-speech gesture synthesis. EMAGE [31] leverages a masked gesture reconstruction to enhance audio-conditioned gesture generation.
4.5 Evaluation Results
@@
Figure 3: Visualization results between the proposed method and other state-of-the-art methods.
Qualitative Results For this task, there is lack of absolutely objective metric to evaluate the performance of various methods. Therefore, We strongly recommend the readers to view our demonstration video to gain an intuitive understanding of the qualitative outcomes. In Figure 3, it shows some synthesized results of the baselines conditioned on the same speech. From these results, we can observe that the synthesized co-speech gestures of the proposed method are more realistic, agile, and diverse than those of the baselines on BEAT datasets. The baselines suffer different extents of jittering in the results, especially lacking of semantic gesture and naturalness. DiffuseStyleGesture [44] is also a diffusion-based method, which directly learn the origin motion representations. However,its results lack smoothness and naturalness (e.g. foot sliding). QPGesture [45] exhibits noticeably slower and less varied motion, while ours demonstrates agility comparable to actual movement. For the specific semantic expressions, apart from our approach, other methods cannot effectively synthesize proper semantic gestures. More results are shown in Figure 4. The results indicate that our method can generate proper semantic gestures for the speech semantics. For example, "free will" refers to the first row with open arms; "left" and "right" indicate pointing to the corresponding direction. These results show the surpassing capability of SIGGesture in synthesizing semantic accurate and natural co-speech gesture.

### v2

@@ 3.2. Semantic Gesture Collection
Figure 2. The statistical data of semantic gesture dataset (upper left part), some examples of semantic label (lower left part), and some semantic gesture examples (right part).
Through extensive observation of numerous instances, it has been discerned that semantic gestures maintain a universality in gestural communication. Specifically, it is evident that different individuals display similar gestural expressions when articulating identical semantic concepts (Lascarides and Stone, 2009; Zhang et al., 2024a). This observation provides a feasible methodology for the collection of an extensive array of semantic gesture actions, aimed at covering a comprehensive range of semantic contexts. We collect a large-scale semantic gesture dataset that contains motion clips $G=\{G_{0},G_{1},...,G_{N}\}$ and corresponding text $T=\{T_{0},T_{1},...,T_{N}\}$ by manual selection and modification with the aid of technical artist. Some of these semantic gestures are borrowed from BEAT dataset. All of the semantic motion clips are shorter than 3 seconds and each motion clip is a whole independent semantic motion. After manual selection and check, we have collected 2,537 semantic motion clips. Some examples of semantic gestures and the distribution statistics are shown in Figure 2. More examples can be found in the supplementary video. These motion clips are sorted into a predefined motion semantic system.
@@
All of semantic gesture clips are encoded into latent embedding $V_{semantic}=\{v_{0},v_{1},...,v_{N}\}$ by a pre-trained VQVAE model. To improve the diversity of semantic gesture, we add a random noise to the embeddings during semantic injection.
3.3. Diffusion models with semantic injection
@@
SIGGesture employs a diffusion model to synthesize the latent representation of gesture, which consists of a diffusion process and a denoising process, as shown in upper part of Figure 1. Given a specific real motion data distribution $p(x_{0})$, our goal is learn a distribution $q_{\theta}(x_{0})$ that is parameterized by ${\theta}$ to approximate the real distribution.
3.3.1. Diffusion process.
@@
We follow the previous DDPM (Ho et al., 2020) definition of diffusion as a Markov noising process with latent $x_{0:T}$ that follows a forward noising process $p(x_{0:T}|x_{0})$. Here, $x_{0}$ is the latent embedding of real motion data. The forward diffusion process is defined as a Markov chain that gradually adds Gaussian noise to the latent representation of real motion, as follows
(1) $\begin{split}q(x_{0:T}|x_{0})=\prod_{t=1}^{T}q(x_{t}|x_{t-1}),\end{split}$
@@
(2) $\begin{split}q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}{I}).\end{split}$
$\beta_{t}$ are variance schedule hyperparameters that control the distribution of Gaussian noise. In the reverse denoising process, $\beta_{t}$ are pre-defined constant parameters. In the diffusion process, the original latent representation of motion $x_{0}$ progressively substituted by random noises. When $T$ approaches infinity, the distribution of $x_{T}$ follows a pure white noise distribution.
3.3.2. Denoising process.
Through diffusion process, the real data are transferred into pure white noise by adding random noise. On the contrary, the denoising process recovers the real data from pure white noise by removing the random noise. The unconditional reverse process of recover a gesture motion embedding can be represented by the following formula
(3) $\begin{split}p_{\theta}(x_{0:T})=p_{\theta}(x_{0})\prod_{t=1}^{T}p_{\theta}(x_{t-1}|x_{t}),\end{split}$
@@
(4) $\begin{split}p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t};u_{\theta}(x_{t},t),\beta_{t}I).\end{split}$
The intermediate state $x_{t}$ is sampled from the following defined distribution
(5) $\begin{split}p(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha_{t}}},(1-\bar{\alpha_{t}}){I}).\end{split}$
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha_{t}}=\prod_{k=0}^{t}\alpha_{k}$. The above explanation shows the process of unconditioned diffusion generation. For conditional content synthesis, the conditions should be injected into the diffusion generation procedure. The details of the denoising network are shown in the upper right part of Figure 1. In our method, the conditions contains audio $c_{audio}$ and speaker identity $c_{id}$. For simple notation, we mark the two conditions as $c=[c_{audio},c_{id}]$. It should be noted that the audio feature is extracted by Jukebox (Dhariwal et al., 2020) instead of WavLM (Chen et al., 2022). Jukebox (Dhariwal et al., 2020), trained on large-scale music datasets, can focus on the rhythmic features rather than the speech semantic transcripts. By injecting the condition $c$ into generation, the reverse process of each time step $t$ can be updated by the following formula
(6) $\begin{split}p_{\theta}(x_{t-1}|x_{t},c)=\mathcal{N}(x_{t};u_{\theta}(x_{t},t,c),\beta_{t}I).\end{split}$
@@
For a speech condition $c$, we reverse the forward diffusion process by learning to estimate $p(x_{t},t,c)$ approximate $x_{0}$ with a parameterized model for all times step $t$. We use a loss function proposed in (Ho et al., 2020) to optimize the whole model, as following,
(7) $\begin{split}\mathcal{L}={E_{x,t}[\parallel x-p(x_{t},t,c)\parallel^{2}_{2}]}.\end{split}$
@@
During training, the model is trained under conditional and unconditional manner with a specific probability. Once the training converges, the diffusion model can predict the noise parameters by considering the unconditional setting and condition $c$.
3.3.4. Inference with semantic injection.
@@
The inference process involves solely the denoising process of diffusion models. The pipeline of inference process is shown in the lower part of Figure 1. To be specific, a speech with its corresponding text will be fed into the generation pipeline. It should be noted that the words in the text are aligned with speech in the timeline. Each word has a duration time. Then, the texts are fed into the LLMs to produce candidate semantic gestures $G_{candicate}=(g_{0},...,g_{N})$ with a pre-defined prompt. $N$ is the number of semantic gestures in this text. The interaction with LLMs process is shown in the left-top corner of inference pipeline in Figure 1. With the ability of LLMs, the proposed method can process various languages, such as Chinese and Japanese.
@@
The candidate semantic gestures have been encoded into latent embedding. We set a control $K$ to determine the degree of semantic injection. From time step $T$ to $K$, the denoising latent variable is marked as $\hat{x_{t}}$ that contains the semantic information.
(8) $\begin{split}p_{\theta}(x_{K:T})=p_{\theta}(x_{T})\prod_{t=K}^{T}p_{\theta}(x_{t-1}|\hat{x}_{t}).\end{split}$
@@
The semantic gesture $G_{candidate}$ can be injected into the latent variable $x_{t}$ by the follow formula
(9) $\begin{split}\hat{x}_{t}=x_{t}*m+G_{candidate}*(1-m),\end{split}$
@@
where $m$ denotes the timeline mask of semantic words. From time step $K$ to 1, the denoising latent variable $x_{t}$ does not contain semantic injection step, which can preserve the structure and smoothness of the generated gesture and improve the diversity of the results.
3.3.5. Generating long sequence.
@@
For synthesizing long motion sequence, the common practice uses auto-regressive manner, which generates content sequentially. However, generative models trained on small pieces may cause unnaturalness and defectiveness. Instead of auto-regressive manner, we adopt DiffCollage (Zhang et al., 2023b) to generate long motion sequence. DiffCollage is a scalable probabilistic model that can synthesize large content in parallel. Specifically, we denote a long sequence as ${\boldmath{m}=[m^{0},m^{1},m^{2}]}$, where $m^{2}$ is the out-painted sequence generated by the conditional model $m^{2}|m^{1}$. Notably, this procedure makes a conditional independence assumption that $q(m^{2}|m^{0},m^{1})=q(m^{2}|m^{1})$.
(10) $\begin{split}q(m)=q(m^{0},m^{1},m^{2})=q(m^{0},m^{1})q(m^{2}|m^{1}).\end{split}$
Further, we can obtain the following formula,
(11) $\begin{split}q(m)=\frac{q(m^{2},m^{1})q(m^{1},m^{0})}{q(m^{1})}.\end{split}$
The score function of $q(m)$ in diffusion model can be represented as a sum over the scores of shorter gesture motions.
(12) $\begin{split}\nabla q(m)=\nabla q(m^{2},m^{1})+\nabla q(m^{1},m^{0})-\nabla q(m^{1}).\end{split}$
From formula 12, it can be found that we can synthesize different gesture clips in parallel since all individual scores can be calculated independently. Then, then we can merge the scores to compute the score of the target long audio condition. For more technical details, please refer to DiffCollage (Zhang et al., 2023b).
3.4. Large-Scale Motion Corpus Collection
@@
In 3D animation, high-quality data collection is remarkably expensive, professional, and time-consuming. High-quality motion capture requires specialized cameras, suits, and markers. Besides, it often requires experienced technicians, actors, and animators to set up, perform, and process the captured data. The process of capturing, cleaning, and integrating motion data is labor-intensive and time-consuming, contributing to the overall expense. In recent years, several high-quality datasets are released to the research community, as shown in Table 1. BEAT (Liu et al., 2022b) is the largest dataset for co-speech gesture synthesis, which contains 76 hours data. However, data-driven approaches with larger models are also short for data. In the inspiration of large-language training paradigm, noise data for pre-training and high-quality data for fine-tuning, we aim at collecting a large-scale gesture corpus for pre-training by video motion capture technique such as HybridIK (Li et al., 2021a), SMPLify (Bogo et al., 2016). To this end, we collect a large mount of videos about speech or talk from TED  [^1] https://www.ted.com and YiXi  [^2] https://yixi.tv. Then, we conduct the data post-processing pipeline to obtain the final shots of interests. Table 2 shows the statistical data of the motion corpus collection. Finally, we obtain about 400 hours 3D gesture assets of speech. As shown in Table 1, the proposed dataset is significantly larger than previous datasets by at least 6 times.
Table 2. Statistical data of the collected motion corpus collection.
@@
Figure 3. Visualization results between the proposed method and other state-of-the-art methods.
4. Experiments
4.1. Datasets
In this paper, we use the largest high-quality speech-gesture dataset BEAT to evaluate the proposed method. BEAT is constructed by using commercial motion capture system, including face emotion and body motion. It contains about 76 hours motion-audio paired data of 30 speakers talking about different topics. We thoroughly post-process the dataset by removing some defective motion sequences such as speaker drinking, wandering, and incorrect motion. Then, the processed BEAT dataset contains 46 hours motion data and its corresponding text and audio. During training the VQVAE model, the original Euler angles are converted to rotation matrix for better convergence. During training the diffusion-based generation model, all motions are represented by latent embeddings in VQVAE latent space. The datasets collected from Internet by video motion capture are only used for pre-training the diffusion model because of its relatively low quality.
4.2. Evaluation Metrics
For content generation tasks, human objective evaluation is the most important evaluation method because of lacking clear ground and explicit criteria. Therefore, the human objective evaluation is the main evaluation method in our experiment. Each motion slice is re-targeted to Mixamo [^3] https://www.mixamo.com model and rendered as a video. The evaluators are asked to rate the slices from the following four aspects respectively: (1) Naturalness (Nat.), (2) Rhythm (Rhy.), (3) Diversity (Div.), (4) Semantic (Sem.). The scores assigned to each rating are in the range of 1-10, corresponding from worst to best. All scores are normalized to 0-1. Besides, we also calculate some quantitative metrics adopted by previous works. FGD (Yoon et al., 2020) measures the distribution difference between generated data and ground truth. Beat Consistency (BC) (Li et al., 2021b) calculates the average distance between every audio beat and its nearest motion beat. Diversity is a metric of measuring the variations of the synthesized gesture. Semantic-Relevant Gesture Recall (SRGR) (Liu et al., 2022b) measures the semantic correct key point of synthesized data by comparing it with the ground truth.
4.3. Implementation Details
The whole training process consists of two steps. Firstly, we train the VQVAE model by using the Adam optimizer with learning rate 0.001 for 400 epochs. Then, we train the diffusion-based gesture generation model using the Adam optimizer with learning rate 0.0002 for 3000 epochs with batch-size 256. The number of diffusion steps for training and inference is set as 1000. During the training of diffusion model, we firstly use the collected data for pre-training (2000 epochs with batch-size 256), and then use the high-quality dataset for fine-tuning (1000 epochs). In the diffusion model, the variances $\beta_{t}$ are predefined to increase linearly from 0.0001 to 0.02, with a total of 1000 noising steps set as $T=1000$. The length of the input audio feature is 12 seconds. We train the diffusion model on 32 V100 GPUs for about one week.
4.4. Comparing Baselines
We compare our method on BEAT dataset with several representative state-of-the-art methods in the last year, including DiffuseStyleGesture (Yang et al., 2023b), TalkSHOW (Yi et al., 2023), QPGesture (Yang et al., 2023a), EMAGE (Liu et al., 2023). DiffuseStyleGesture (Yang et al., 2023b) is a diffusion model-based co-speech gesture synthesis approach, which takes the motion style into generation process. TalkSHOW (Yi et al., 2023) uses a simple full-body speech-to-motion generation framework with auto-regressive manner, in which the face, body, and hands are modeled separately. QPGesture (Yang et al., 2023a) is a quantization-based and phase-guided motion-matching approach for co-speech gesture synthesis. EMAGE (Liu et al., 2023) leverages a masked gesture reconstruction to enhance audio-conditioned gesture generation.